Harry potter spend 2/3 of his allowance at the movies night. his mom was to sweet and let him earn 12 additional dollars for washing the car on Saturday. if Harry potter ended the week with $19,
Answer by fractalier(6550)

Call what he started with, x.
Then x - (2/3)x + 12 = 19
(1/3)x + 12 = 19
(1/3)x = 7
x = $21